OpenStudy (anonymous):

Find the derivative of f(x) = 4/x at x = 2

OpenStudy (danjs):

power rule 4x^(-1)

OpenStudy (danjs):

f ' (x) = -4x^(-2)

OpenStudy (danjs):

put in x = 2

OpenStudy (anonymous):

-4(2)^(-2) -4(.25) -1?

OpenStudy (anonymous):

is -1 correct?

OpenStudy (danjs):

-4/x^2 = -4/4 = -1

OpenStudy (danjs):

sorry i misstyped the first one

OpenStudy (danjs):

any other probs?

OpenStudy (anonymous):

Thats it for now! Thank you!